Grade 11Physical Chemistry

the heat of combustion of Ethanol determined in a Bomb calorimeter is -670.48 km in per kg per mole at 25 degree Celsius what is delta Hat 25 degree Celsius for the reaction

Zaid
7 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago
similar question
Let us first write reaction of combustion in benzene.

C6H6 + 7½O2 (g) -----> 6CO2 (g) + 3H2O (l)

In the above reaction O2 and CO2 are the only gases.
g = no. of moles of products – no. of moles of reactants
= 6 - 7½ = - 1 ½ = -3/2

Given valuesU = 3263.9 kJ/mol; T = 298 K;
As we know,
H =U +ng RT R = 8.314

As this is an exothermic reactionH is negative.

H = (- 3263.9 kJ/mol) + (-3/2) (298 K)
= (- 3263.9 kJ/mol) + (-3/2) (2.477)
= (- 3263.9 kJ/mol) – 3.716 kJ/mol
H = - 3267.6 kJ/mol